# Distilling heterogeneous treatment effects: Stable subgroup estimation in causal inference

### v1

@@ 5 Simulations
In the following section, we demonstrate the effectiveness of CDT in accurately estimating subgroups through extensive simulations. We consider two different teacher models: causal forest (Wager and Athey, 2018) and R-learner with boosting (rboost) (Nie and Wager, 2021). We benchmark CDT’s performance, relative to popular subgroup estimation approaches — namely, causal trees (Athey and Imbens, 2016), virtual twins (Foster et al., 2011), and interacted linear models, fitted without regularization (i.e., ordinary least squares) and with Lasso ($L_{1}$) regularization (e.g., Imai and Ratkovic, 2013). Details regarding the implementation of each CDT and comparison method can be found in Appendix C. Across the different simulation scenarios, we find that CDT consistently estimates the true subgroup features, thresholds, and CATEs with greater accuracy than existing methods. Moreover, our model selection procedure using the Jaccard SSI provides an effective data-driven approach to choosing an appropriate teacher model for CDT.
5.1 Simulation Setup
We consider a random treatment assignment process, where $Z_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Bernoulli}(1/2)$ for $i=1,\ldots,n$ samples. Furthermore, we generate $n$ samples and $p$ covariates from a standard multivariate normal distribution (i.e., $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{MVN}(0,I)$, where $X_{i}\in\mathbbm{R}^{p}$ and $i=1,\ldots,n$).
To capture a broad range of subgroup data-generating processes (DGPs), we consider three different treatment effect heterogeneity models:
@@ 5.1 Simulation Setup
where $\epsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{\tau}^{2})$. The goal of the simulation study is to evaluate whether the different subgroup estimation methods can accurately find the subgroups associated with $\tau(X_{i})$. For example, for the ‘AND’ subgroup DGP, an accurate subgroup estimation method would correctly estimate $\mathbbm{1}\left\{X_{i}^{(1)}>0\right\}\cdot\mathbbm{1}\left\{X_{i}^{(2)}>0.5\right\}$ as a subgroup. For each setting, we also vary $\sigma_{\tau}$ such that the proportion of variance explained in $\tau(X)$ by the covariates $X$ vary across the values $\{0.2,0.4,0.6,0.8,1\}$. [^4] The proportion of variance explained in $\tau(X)$ by $X$ is defined as $\text{var}(\mathbb{E}[\tau_{i}\mid X])/\text{var}(\tau_{i})\in[0,1]$.
The outcome process is a function of the underlying treatment effect heterogeneity model:
$Y_{i}=Z_{i}\cdot\tau_{i}+X_{i}^{(3)}+X_{i}^{(4)}+\nu_{i},$
where $\nu_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,0.1^{2})$ is an additive noise term.
  1. 1.
  2. 2.
  3. 3.
For each simulation replicate, we generate $n=500$ samples and $p=10$ features. We consider addition data generating processes in Appendix C, but find that the results are largely consistent, even across more complex settings.
5.2 Main Simulation Results
Across the different simulation settings, we find that the CDT methods are able to accurately estimate the true subgroups, both in terms of finding the correct features (as measured by $F_{1}$ score), as well as estimating the correct thresholds to construct the decision rules. In contrast, standard subgroup estimation methods (i.e., causal trees and virtual twins) result in many erroneously constructed subgroups, often overfitting to noise features, while linear approaches (i.e., linear regression and Lasso) often fail to pick up on existing subgroups. Figure 4 visualizes the key simulation results, with more detailed summaries in Appendix C.
There are several key takeaways from the simulation results. First, the simulations highlight the advantages of distillation for subgroup estimation. In particular, when comparing distillation with traditional tree-based subgroup estimation methods (i.e., causal tree and virtual twins), we find that the relative improvement in $F_{1}$ score from distillation arises from reducing the number of false positives. In particular, virtual twins and causal tree often split on noise features, particularly, $X^{(3)}$ and $X^{(4)}$, while the distilled methods avoid selecting such features (see Appendix C). In other words, without distilling, the tree-based approaches end up erroneously constructing subgroups. This reinforces our theoretical findings from Section 3 that distillation can improve the stability in recovering subgroups, and mitigate potential overfitting in the presence of noise. In comparing the estimated subgroup thresholds for $X^{(1)}$ and $X^{(2)}$, we find that the CDT methods also yield the most accurate and precise estimates of the subgroup thresholds for $X^{(1)}$ and $X^{(2)}$ across the various subgroup DGPs and treatment effect heterogeneity strengths.
We can similarly compare distillation with linear-based approaches (i.e., linear regression and Lasso). In higher noise settings, the linear approaches suffer from the opposite problem as standard tree-based methods. Instead of overfitting to create erroneous subgroups, the linear methods fail to detect any subgroups. This results in substantially lower true positive rates when compared to CDT. In contrast, CDT is still able to detect subgroups in high noise settings. When moving to lower noise settings (i.e., DGPs with moderate-to-high signal strengths), we see that CDT generally outperforms the linear approaches on every accuracy measure (i.e., lower false positives, higher true positives, and higher $F_{1}$ scores).
Second, the simulations highlight advantages from distillation in estimating the subgroup ATEs. In particular, improvements in accurately recovering the subgroups result in lower RMSE in recovering the subgroup ATEs. Moreover, we find that the performance of the traditional tree-based approaches and the linear methods depend heavily on the structure of the underlying outcome model. For example, in the DGP considered in the main manuscript, the outcome model depends on linear functions of the other covariates (namely, $X^{(3)}$ and $X^{(4)}$). Here, the linear-based methods, unsurprisingly, perform well, while the traditional tree-based approaches frequently perform worse in noisy regimes. In contrast, in Appendix D, we additionally consider settings when the outcome model is soley a function of the conditional average treatment effect. In these settings, the existing tree-based methods yield relatively low subgroup ATE errors, while the linear methods result in substantially higher subgroup ATE RMSE. In comparison, the distillation methods appear to be relatively robust to the underlying outcome model’s specification, and perform well in both settings.
@@ 5.2 Main Simulation Results
Figure 4: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via root mean squared error (RMSE) for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes with linear covariate effects (rows). CDT with various teacher models (i.e., Distilled Causal Forest and Distilled Rboost) frequently yields the highest $F_{1}$ and number of true positives alongside the lowest number of false positives, threshold RMSEs, and CATE RMSE, demonstrating its effectiveness for accurate subgroup estimation. Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
5.3 Teacher Model Selection Simulations
@@
In addition to the subgroup estimation performance, we also examine our proposed model selection procedure using the Jaccard SSI under each of the aforementioned simulation scenarios. We compute the Jaccard SSI for both Distilled Causal Forests and Distilled Rboost using $B=100$ bootstrap samples and four different choices of tree depths ($d=1,2,3,4$). As a baseline, we also compute the Jaccard SSI for the causal tree. Figure 5 provides a visual summary of the results.
@@
We find that regardless of the noise level, Distilled Causal Forest almost uniformly yields the highest subgroup Jaccard stability score, even varying different tree depths. This implies that according to our proposed teacher model selection procedure, researchers should be using a Causal Forest instead of Rboost as the teacher model. We can directly check the performance of the two teacher models, and find that indeed, the Distilled Causal Forest resulted in higher $F_{1}$ scores for the selected subgroup features, and a lower CATE RMSE as compared to Distilled Rboost.
In general, we see that a higher Jaccard SSI generally corresponds to more accurate subgroup estimation. A similar pattern emerges when examining other simulation scenarios (see Appendix E for additional discussion), thereby demonstrating the efficacy of our stability-driven procedure to select the teacher model.
@@ 6 Case Study: AIDS Clinical Trials Group Study 175
To illustrate the stability and interpretability of CDT on real data, we turn to the AIDS Clinical Trials Group Study 175 (also referred to as ACTG 175). ACTG 175 was a randomized controlled trial to determine the effectiveness of monotherapy compared to combination therapy on HIV-1 infected patients. To enter the study, participants’ CD4 cell counts had to be between 200 to 500 cells per cubic millimeter upon screening. The main outcome of interest was whether a patient reached the primary end point ($Y_{i}=0$), defined by at least a 50% decline in CD4 cell count, the development of acquired immunodeficiency syndrome (AIDS), or death.
[…]
Subgroup Results
[…]
Combination therapy was shown to slow the progression of HIV-1. However, our goal in this case study is to further characterize subgroups of the population that may experience varying effectiveness of combination therapy. Although we have no underlying ground truth to compare results with as in our simulation study, given substantial medical advances and follow-up research has been performed since this study was conducted in 1996, we can at least compare to current knowledge to vet our results. Here, we focus specifically on the impact of a combination of zidovudine and zalcitabine as our treatment compared to a monotherapy regime that only includes zidovudine. We used a combination of medical records and demographic information as covariates to estimate subgroups and their corresponding treatment effects.
6.1 Estimating subgroups
To illustrate a standard application of the proposed method, we apply CDT using Rboost as the teacher model to a 50% split of the study and estimate subgroup average treatment effects on the remaining 50%. Note we are focusing on results from Distilled Rboost because it yielded a higher subgroup Jaccard stability score across varying tree depths (see Appendix F). In Table 1, we provide the estimated treatment effects and descriptions of each subgroup, and in Figure 6, we visualize the resulting tree structure, where values in the leaf nodes denote the estimated subgroup effects. Across estimated groups, the average number of treatment and control samples is 32 and 34 respectively.
Interactions between CD8 cell counts, weight at the start of the study, and the amount of previous exposure to anti-retroviral therapy are the key characeristics forming our estimated subgroups. Interestingly, these variables are now well-known to influence the progression of HIV-1. For instance, the HIV-1 viral load and resulting decrease in CD4 cell counts traditionally was the major marker to measure response to treatment. However, absolute CD4 counts may not accurately reflect the risks facing patients given immune dysfunction persists despite normalization of counts (McBride and Striker, 2017). The CD4/CD8 ratio, where a value above 1 is considered healthy, is now known to more accurately characterize overall immune dysfunction and may be a stronger indicator of disease progression, response to treatment, morbidity, and mortality. For added context on our estimated subgroups, in this study 99% of individuals had CD4 cell counts below 700, following the rule of thumb that a healthy CD4/CD8 ratio is above 1, CD8 values above 700 would place the majority in an unhealthy range.
Similarly, zidovudine is associated with lipodystrophy (Finkelstein et al., 2015)—a condition that affects how the body stores fat resulting in weight gain in harmful parts of the body (e.g., organs) and loss in others. Thus, a higher or much lower starting weight can put individuals at higher risk of reaching the primary endpoint. Finally, those with previous anti-retroviral therapy improve their chances of longer life expectancy and balanced CD4/CD8 counts. Thus, those entering the study with less experience may benefit the most from a more aggressive treatment regime.
These three covariates clearly interact in specific ways across our eight subgroups. For instance, for those with high CD8 counts, combination therapy is more effective when individuals also have a higher starting weight and less treatment exposure as in Group 6. The effectiveness decreases when combined with a lower starting weight (Group 5). Even if CD8 counts are lower (i.e., bringing the CD4/CD8 ratio closer to 1), a similar though slightly weaker effect is observed with a stronger impact on those with less treatment exposure and a higher starting weight (Group 2). In contrast, a lower starting weight and less treatment hinders the effectiveness (Group 1). Note while in Table 1, we are referring to “low to medium” CD8 counts, this is relative to our study population. As discussed above, this includes CD8 counts well above 700, placing most in the unhealthy range.
@@ 6.1 Estimating subgroups
To summarize, CDT recovers clinically-relevant subgroups in the AIDS Clinical Trial data that we now know to be substantively important, thus providing a helpful validity check on the usefulness of the proposed method. However, this illustration also highlights that had researchers utilized CDT when the initial trial was conducted, they could have identified these potential drivers of treatment effect heterogeneity, without having to a priori specify what to search for.
6.2 Stability analysis
To consider the stability of the different methods, we perform a bootstrap analysis. Across 100 different bootstrap samples, we apply CDT, virtual twins, as well as tree-based methods and linear methods, to compare the different subgroups obtained under each technique. For each bootstrap instance, we compute the proportion of times features were used to construct subgroups. We summarize the results across these runs in Figure 6.
Figure 6: For our case study, we summarize (A) the resulting decision tree produced by CDT using a Causal Forest as the teacher model and (B) the proportion of times across the 100 bootstrap instances features are split upon to form rules and compose subgroups across methods (colors). Proportions closer to one indicate more stability and robustness to small data perturbations.
Comparing the CDT results with the causal tree, we see that the causal tree also similarly constructs subgroups using CD8 cell count, participants’ starting weight, and the number of days of pre-175 antiretroviral therapy. However, it additionally splits on other covariates, depending on the bootstrap iteration, though with less consistency. Virtual twins consistently constructs subgroups on those same three features, but also constructs substantially more across several of the other features producing more subgroups that are challenging to interpret.
Interestingly, the linear-based approaches (i.e., linear regression and Lasso) generally fail to pick up on any of the features, except for the Karnofsky score. The Karnofsky score is an assessment tool intended to assist clinicians and caregivers in gauging a patient’s functional status and ability to carry out activities of daily living (Schag et al., 1984). While it would normally be assumed to be an informative and clinically relevant feature, the study eligibility criteria required that all participants have a performance score of at least 70; in fact, the median score in the study was 100, translating to a normal status with no evidence of disease. As a result, there is little heterogeneity across the Karnofsky score of the patients in the study and it would not likely moderate the treatment in a meaningful way.
@@ 7 Discussion
In the following paper, we introduced a novel method, causal distillation trees, for interpretable subgroup estimation in causal inference. CDT allows researchers to leverage the flexibility and power of black-box metalearner approaches, while preserving the interpretability of simple decision trees. We prove that CDT allows researchers to consistently estimate the optimal set of subgroups, and we introduce a novel model selection procedure to help researchers use CDT in practice. To help researchers implement CDT in practice, we have developed an R package causalDT available at https://github.com/tiffanymtang/causalDT, which not only fits CDT but also automatically conducts the subgroup stability and diagnostic checks, thus enabling researchers to interpret their estimated subgroups from CDT alongside important and informative diagnostics in a holistic manner.
There are several promising avenues for future research. First, recent literature in external validity has emphasized the importance of recruiting a sufficiently heterogenous experimental sample that adequately captures the total treatment effect variation across the target population (e.g., Huang and Parikh, 2024; Egami and Lee, 2024; Hoellerbauer and Laterzo-Tingley, 2024; Tipton and Mamakos, 2023; Zhang et al., 2024). Failing to do so can result in systematically omitting key segments of the target population, making it infeasible to generalize the causal effect from an experimental sample to the target population of interest (e.g., Huang, 2024). CDT offers an opportunity for researchers to characterize the underlying treatment effect heterogeneity in a given study and to consider how to optimally recruit these subgroups of participants for an externally valid study. Integrating CDT into external validity concerns would be an interesting avenue of future research.
Second, CDT provides a very general and flexible two-stage framework. Throughout the paper, we focused on CART as the second-stage learner. However, there are numerous alternative tree-based models (Loh, 2002; Quinlan, 2014; Hu et al., 2019) and many that have been developed specifically for subgroup estimation (e.g., Loh et al., 2015; Su et al., 2009; Lipkovich et al., 2011; Seibold et al., 2016; Wang et al., 2024), which can be used as second-stage learners in CDT and may further improve performance. Theoretical properties of these learners remain understudied. Similarly, efforts to further enhance the stability of distillation-based methods (Zhou et al., 2024) can be easily incorporated into the broader CDT framework.
Finally, while the paper considers the causal inference setting, where researchers are interested in estimating relevant subgroups across which they expect there to be treatment effect heterogeneity, an extension of CDT could consider prediction models, where there may be variation in the underlying predictive performance of a specific model. Thoroughly vetting models based on predictive stability has improved the interpretability and predictive performance across several substantive areas including cancer screening and genetic drivers of heart disease (Irajizad et al., 2023; Wang et al., 2023), and a more in-depth approach at the subgroup level could help diagnose potential disparities in the performance of predictive models. Furthermore, while splits from our second stage learner were directly used to characterize subgroups, substantial work has been done to improve feature importances on black-box methods (Lundberg, 2017), especially tree-based approaches (Agarwal et al., 2023; Li et al., 2019), to determine how heavily different features influence predictive performance. These approaches can be combined with CDT to inform which characteristics will most heavily impact the subgroup estimation.

### v2

@@ 5 Simulations
In this section, we demonstrate the effectiveness of CDT in accurately estimating subgroups through extensive simulations. We consider three different teacher models: R-learner with boosting (Rboost), causal forest, and Bayesian causal forest (BCF). We benchmark CDT’s performance, relative to popular subgroup estimation approaches — namely, causal trees (Athey & Imbens, 2016), virtual twins (Foster et al., 2011), and interacted linear models, fitted without regularization (i.e., ordinary least squares) and with Lasso ($L_{1}$) regularization (e.g., Imai & Ratkovic, 2013). Implementation and method details can be found in Appendix C. Across the different simulation scenarios, we find that CDT consistently estimates the true subgroup features, thresholds, and CATEs with greater accuracy than existing methods. Moreover, our model selection procedure using the Jaccard SSI provides an effective data-driven approach to choosing an appropriate teacher model for CDT.
Simulation Setup.
We generate a random treatment assignment process, where $Z_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{Bernoulli}(1/2)$), and generate $n=500$ samples and $p=10$ covariates from a standard multivariate normal distribution (i.e., $X_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\text{MVN}(0,I)$, where $X_{i}\in\mathbbm{R}^{p}$ and $i=1,\ldots,n$). Furthermore, we simulate three different treatment effect heterogeneity models:
@@
where $\epsilon_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\sigma_{\tau}^{2})$, and we define the outcome process to be a function of this underlying treatment effect heterogeneity: $Y_{i}=Z_{i}\cdot\tau_{i}+X_{i}^{(3)}+X_{i}^{(4)}+\nu_{i},$ where $\nu_{i}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,0.1^{2})$. We vary $\sigma_{\tau}$ such that the proportion of variance explained in $\tau(X)$ by the covariates $X$ vary across the values $\{0.2,0.4,0.6,0.8,1\}$. [^4] The proportion of variance explained in $\tau(X)$ by $X$ is defined as $\text{var}(\mathbb{E}[\tau_{i}\mid X])/\text{var}(\tau_{i})\in[0,1]$. To evaluate the effectiveness of each subgroup estimation method, we consider how accurately each method recovers (1) selected subgroup features, (2) estimated subgroup thresholds, and (3) estimated subgroup ATEs (details in Appendix D). We also consider additional data generating processes in Appendix D, and find that the results are largely consistent, even across more complex settings.
@@
Figure 3: Performance of subgroup estimation methods for (A) identifying the true subgroup features, measured via $F_{1}$ score, number of true positives, and number of false positives, (B) estimating the true subgroup thresholds, measured via RMSE for each true subgroup feature, and (C) estimating the true subgroup ATE, measured via RMSE, across increasing treatment effect heterogeneity strengths (x-axis) and different subgroup data-generating processes with linear covariate effects (rows). Results are averaged across 100 simulation replicates with ribbons denoting ±$\pm 1SE$.
Subgroup Estimation Results.
Figure 3 visualizes the key simulation results, with more detailed summaries in Appendix D. Across the different simulation settings, we find that the CDT methods are able to accurately estimate the true subgroups in terms of finding the correct features (as measured by $F_{1}$ score in Figure 3A), estimating the correct thresholds to construct the decision rules (Figure 3B), and estimating the true subgroup ATEs (Figure 3C). In contrast, standard subgroup estimation methods (i.e., causal trees and virtual twins) often result in many erroneously constructed subgroups, overfitting to noise features (shown by the high number of false positives). This reinforces our theoretical findings from Section 3 that distillation helps mitigate potential overfitting in the presence of noise. Meanwhile, linear approaches (i.e., linear regression and Lasso) often fail to pick up on existing subgroups (shown by the low number of true positives). The simulations further highlight advantages from distillation in estimating the subgroup ATEs. In addition to generally yielding the lowest RMSE in recovering the subgroup ATEs, we find that CDT is relatively robust to the underlying outcome model’s specification while some existing methods, namely, the linear methods, are highly sensitive to the form of the outcome model (see Figure A3 and Appendix D for details).
Teacher Model Selection Simulations.
@@
To examine our proposed teacher model selection procedure, we computed the Jaccard SSI for Distilled Rboost, Distilled Causal Forest, and Distilled BCF using $B=100$ bootstrap samples and four different choices of tree depths ($d=1,2,3,4$). As a baseline, we also computed the Jaccard SSI for the causal tree.
@@
Figure 4 provides a visual summary of the results under the ‘AND’ subgroup simulation. Under this simulation, we find that regardless of the noise level, Distilled Causal Forest almost uniformly yields the highest subgroup Jaccard stability score, even varying different tree depths. According to our proposed teacher model selection procedure, researchers should thus be using Causal Forest instead of Rboost or BCF as the teacher model. We can directly check the performance of the two teacher models, and find that indeed a higher Jaccard SSI generally corresponds to more accurate subgroup estimation. A similar pattern emerges when examining other simulation scenarios (see Appendix E), thereby demonstrating the efficacy of our stability-driven procedure to select the teacher model.
@@ 6 Case Study: AIDS Clinical Trials Group Study 175
To illustrate the stability and interpretability of CDT on real data, we turn to the AIDS Clinical Trials Group Study 175 (ACTG 175). ACTG 175 was a randomized controlled trial to determine the effectiveness of monotherapy compared to combination therapy on HIV-1 infected patients. To enter the study, participants’ CD4 cell counts had to be between 200 to 500 cells per cubic millimeter upon screening. The main outcome of interest was whether a patient reached the primary end point ($Y_{i}=0$), defined by at least a 50% decline in CD4 cell count, the development of acquired immunodeficiency syndrome (AIDS), or death.
Given that combination therapy is already known to slow the progression of HIV-1, our goal is to further characterize subgroups of the population that may experience varying effectiveness of combination therapy using clinical and demographic variables as pre-treatment covariates. Moreover, we specifically focus on the impact of zidovudine and zalcitabine as our treatment compared to a monotherapy regime using zidovudine.
To begin, we perform the proposed model selection procedure to a 50% split of the study, and find that Rboost results in the highest Jaccard SSI across varying tree depths. We then apply CDT, using Rboost as the teacher model, and estimate subgroup ATEs on the remaining 50% (see Table 1 for the estimated treatment effects and descriptions of each subgroup and Figure 5 for a visualization of the resulting tree).
The key covariates in our estimated subgroups are CD8 cell counts, weight at the start of the study, and the amount of previous exposure to anti-retroviral therapy. Importantly, these variables are well-known to influence the progression of HIV-1. CD8 cell counts in particular are used in combination with CD4 counts (McBride & Striker, 2017) to characterize overall immune dysfunction, and a CD4/CD8 ratio above 1 is considered healthy. [^5] For added context on our estimated subgroups, in this study 99% of individuals had CD4 cell counts below 700, following the rule of thumb that a healthy CD4/CD8 ratio is above 1, CD8 values above 700 would place the majority in an unhealthy range. Zidovudine is associated with lipodystrophy (Finkelstein et al., 2015)—a condition that affects how the body stores fat resulting in weight gain in harmful parts of the body (e.g., organs) and loss in others. Thus, a high starting weight can put individuals at higher risk of reaching the primary endpoint. Finally, while previous anti-retroviral therapy can initially help improve CD4 counts, over time, individuals can develop drug resistance and benefit from a different regime.
These three covariates interact across our subgroups, producing varying treatment effects. Those with a higher starting weight benefit more from combination therapy under both lower and higher CD8 counts (i.e., Groups 2 and 3 vs Group 1), though the effect is slightly stronger under a lower (healthier) CD8 count. Notably, it is only when those with a lower starting weight also have the highly risky combination of high CD8 cell counts and substantial previous therapy 666786 days of therapy is in the upper 78th percentile of previous exposure in this study that they benefit significantly from the treatment (Group 4).
@@
To summarize, CDT recovers clinically-relevant subgroups in the AIDS Clinical Trial data that we now know to be substantively important, thus providing a helpful validity check on the usefulness of the proposed method. This illustration also highlights that had researchers utilized CDT when the initial trial was conducted, they could have identified these potential drivers of treatment effect heterogeneity, without knowing them a priori.
We also evaluate the stability of the different methods across bootstraped samples, and find distillation methods both consistently split on the same set of features and that estimated subgroups were overwhelmingly characterized by features with direct clinical relevance. In contrast, the other tree-based approaches split on a variety of different features, depending on the bootstrapped sample, while the linear approaches frequently failed to identify meaningful heterogeneity. See Appendix F.3 for details.
Figure 5: For our case study, we have the resulting decision tree produced by CDT using Rboost as the teacher model. Estimated subgroup ATEs are shown (SE from 1000 bootstrapped samples in parenthesis) and significant effects are outlined in blue.
@@ 7 Discussion
We introduced causal distillation trees for interpretable subgroup estimation in causal inference. CDT allows researchers to leverage the flexibility and power of black-box metalearner approaches, while preserving the interpretability of simple decision trees. We prove that CDT allows researchers to consistently estimate the optimal set of subgroups, and we introduce a novel model selection procedure to help researchers use CDT in practice. Furthermore, we have developed an R package causalDT available at https://github.com/tiffanymtang/causalDT, which not only fits CDT but also automatically conducts the subgroup stability and diagnostic checks, thus enabling researchers to interpret their estimated subgroups from CDT alongside important and informative diagnostics in a holistic manner.
There are several promising avenues for future research. First, recent literature in external validity has emphasized the importance of recruiting a sufficiently heterogenous experimental sample that adequately captures the total treatment effect variation across the target population (e.g., Huang & Parikh, 2024; Huang, 2024+). CDT offers an opportunity for researchers to characterize the underlying treatment effect heterogeneity in a given study and to consider how to optimally recruit these subgroups of participants for an externally valid study. Second, CDT provides a very general and flexible two-stage framework. While we focused on CART as the second-stage learner, there are numerous alternative tree-based models (e.g., Loh, 2002; Quinlan, 2014; Hu et al., 2019) and many that have been developed specifically for subgroup estimation (e.g., Loh et al., 2015; Su et al., 2009; Lipkovich et al., 2011; Seibold et al., 2016; Wang et al., 2024), which can be used as second-stage learners in CDT and may further improve performance. Similarly, efforts to further enhance the stability of distillation-based methods (Zhou et al., 2024) can be easily incorporated into the broader CDT framework. Lastly, while this work is focused on the causal inference setting and treatment effect heterogeneity, an extension of CDT could consider prediction models, where there may be variation in the underlying predictive performance of a specific model. In this prediction context, subgroup analyses could help diagnose critical disparities in the performance of predictive models.